If the velocity distribution of a fluid over a plate is given by u = (3/4)y – y 2 , where u is the velocity

in metre per second at a distance of y metres above the plate, determine the shear stress at

y = 0.15 metre. Take a dynamic viscosity of a fluid as 8.5 x 10 -5 kg.s/m 2 .


1
Expert's answer
2021-05-31T16:48:29-0400

Gives

Velocity


"u=\frac{3}{4}y+y^2"

y=0.15 meter

"\eta=8.5\times10^{-5}kg s/m^2"

"\frac{du}{dy}=\frac{3}{4}+2y"

At y=0.15meter

"\frac{du}{dy}=\frac{3}{4}+2(0.15)"

"\frac{du}{dy}=0.75+0.3"

"\frac{du}{dy}=1.05"

Shear stress = "\eta\times\frac{du}{dy}"

Put value

Shear stress =


"8.5\times10^{-5}\times1.05=8.925\times10^{-05}N/m^2"


Shear stress="8.925\times10^{-05}N/M^2"